# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ IV Dynamical Equations
IV.1 Dynamical equations: commutative approach
Going a few steps back, the Hořava-Lifshitz commutative BCG action (2) is related to the Lagrangian density ${\cal L}_{HL}$, expressed in terms of the standard variable $u(t)$ in the form [8] (see also refs. [38, 50, 49])
@@ IV.1 Dynamical equations: commutative approach
$p_{u}=\frac{\partial{\cal L}_{HL}}{\partial\dot{u}}=-\frac{u\dot{u}}{N}\,.$ (41)
@@
${\cal H}_{HL}=\frac{1}{2}\frac{N}{u}\Biggl{[}-p^{2}_{u}+g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{]}\,.$ (43)
Figure 4: Graphical illustrations of the potential $V(u)$ of equation (50), emphasizing the contour lines. The values of the running coupling constants and parameters are, in the upper figures, $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$. In the lower figure, the values are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\omega=-1$ ${\cal C}=1$.
@@
${\cal H}_{fp}=\frac{1}{2}N\frac{p_{v}}{u^{3\omega}}\,,\quad\mbox{with}\quad p_{v}=-\frac{v\dot{v}}{N}\,,$ (44)
resulting, from equation (7), assuming the gauge $N=1$, and simplifying the notation, eliminating the HL label, in the super-Hamiltonian:
=12[${\cal H}\!=\!\frac{1}{2}\Biggl{[}-\frac{1}{u}p^{2}_{u}+\frac{g_{r}}{u}-g_{m}-g_{k}u-g_{q}u^{2}+g_{\Lambda}u^{3}+\frac{g_{s}}{u^{3}}+\frac{1}{u^{3\omega}}p_{v}\Biggr{]}.$ (45)
From equation (45) the following Hamilton equations result:
@@
$\dot{p}_{v}$ $=$ $-\frac{\partial{\cal H}}{\partial v}=0\rightarrow p_{v}={\cal C}\,,$ (46)
@@
$\dot{p}_{u}=-\frac{\partial{\cal H}}{\partial u}$ $=$ $-\frac{1}{2u}\Biggl{[}\frac{p^{2}_{u}}{u}-\frac{g_{r}}{u}-g_{k}u-2g_{q}u^{2}\Biggr{]}$ (47)
-- $\frac{1}{2u}\Biggl{[}3g_{\Lambda}u^{3}-3\frac{g_{s}}{u^{3}}-3\omega\frac{{\cal C}}{u^{3\omega}}\Biggr{]}\,,$
where ${\cal C}$ is a constant in the time sector. It is important to point out that the relative sign between the term labeled by $g_{\Lambda}$ has changed, in the derivative procedure, in comparison with the $g_{r}$ and $g_{s}$ contributions (see equation (45)).
IV.1.1 Second-order dynamical equations: commutative formulation
@@
$p_{u}=-u\dot{u}\,,\quad\mbox{and}\quad\dot{p}_{u}=-\dot{u}^{2}-u\ddot{u}\,,$ (48)
we obtain from (47) the following dynamical equation for the scale factor $u$:
$2u\ddot{u}+\dot{u}^{2}+V(u)=0\,.$ (49)
@@
$V(u)=g_{k}+2g_{q}u-3g_{\Lambda}u^{2}+\frac{g_{r}}{u^{2}}+3\frac{g_{s}}{u^{4}}+3\omega\frac{{\cal C}}{u^{3\omega+1}}\,.$ (50)
Figure 4 show typical graphical illustrations of the potential $V(u)$, eq. (50). The main difference between the curves is related to the sign of the stiff matter running coupling constant, for positive signal it gives rise to a repulsive potential (upper curves) and for negative values an attractive potential around a singularity in the origin.
Figure 5: Sample solutions for the scale factor $u(t)$ given by equation (49) for initial conditions $u(0)=0.15$, 0.35, and 1.0. The values of the running coupling constants and parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$.
Figures 5 show typical sample solutions of the second-order time derivative of the scale factor $u(t)$ given in equation (49). For all values of the running coupling constants, the results indicate, differently from generally assumed, that the Universe does not originate from a singularity or out of nothing as originally proposed by Alan Guth [6, 7] in the inflation model (see also ref. [66]). Our results are more in tune with a transition phase between the contraction and expansion topological boundaries intermediated by a Riemannian foliated transition region as proposed by the branch-cut gravity.
@@
Figure 6: Graphical illustrations of the potential $V(u,t)$ and the scale factor of equations (53) and (54) for the commutative approach. In the upper pair of figures, the values of the running coupling constants and parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$. In the below pair of figures, the values are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\omega=-1$ ${\cal C}=1$.
IV.1.2 First-order dynamical equations: commutative formulation
In what follows, we use Cauchy’s implicit function theorem, assuming that the variable $u$ has an explicit time dependence, so the variables $u$ and $t$ can be separated, thus ensuring that these variables are differentiable. We also assume, for simplicity, in view of the high difficulty to solve the above equations, that $\dot{u}$ and $u$ act as independent variables, so from equation (49) we have
$2u\frac{d\dot{u}}{dt}+\dot{u}\frac{du}{dt}=-V(u)\rightarrow 2ud\dot{u}+\dot{u}du=-V(u)dt\,.$ (51)
@@ IV.1.2 First-order dynamical equations: commutative formulation
$V(u,t)\equiv\frac{1}{3}\Biggl{(}\frac{g_{k}}{u}+2g_{q}-3g_{\Lambda}u+\frac{g_{r}}{u^{3}}+3\frac{g_{s}}{u^{5}}+3\omega\frac{{\cal C}}{u^{3\omega+2}}\Biggr{)}t\,.$ (54)
@@
Figure 6 presents graphical representations of equation (54). In these contour plots, varying colors indicate the behavior of the potential $V(u,t)$ in regions of dominance, either repulsion or attraction. Lighter colors signify higher intensity or amplitude, while darker colors indicate the opposite. The blue color signifies the region with lower values, typically surrounding the singularity or the repulsive part of the potential, while the cream color designates regions with higher values, where either the singularity or the repulsion of the potential predominates. The white regions, albeit not always clearly visible in some of the figures, corresponds to the exact region of the singularity or the maximum of repulsion. The results regarding the potential $V(u,t)$ are characterized by symmetric distributions of color compositions, indicating that the commutative algebraic structure contributes in a moderate way to the configuration and distribution of matter in the early Universe.
@@
Figure 7 displays typical sample solutions of equation (53) for various sets of initial conditions. The outcomes obtained from the dynamic equations suggest that, contingent upon the initial conditions, phases that lead to the big crunch or moderately accelerated phases of the early Universe tend to prevail.
@@
Figure 7: Sample solutions of the first-order time derivative of $u(t)$ as given in equation (53) for the commutative formulation. The initial condition is $u(0)$. The values of the running coupling constants and parameters are as follows: Upper-left figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$; Upper-right figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=-0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$; Below-left figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\omega=-1$; ${\cal C}=1$; Below figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=-0.4$; $g_{s}=-0.03$; $\omega=-1$; ${\cal C}=1$.
@@
The consistency of the assumptions that led to equation (53) and to the potential $V(u,t)$ given by equation (54) will be tested in the following by comparing the results that describe the behavior of the solutions of the dynamic equations (49) and (53). It is important to emphasize that articles addressing this topic typically do not provide solutions to the equations describing the behavior of the Universe’s wave function due to the formal challenges associated with these descriptive equations. These approaches are constrained to performing computational or algebraic calculations for dynamic equations, typically relying on approximations that significantly restrict their scope and predictive capabilities. In our case, from a computational perspective, we utilized the Runge-Kutta-Fehlberg method for solving differential equations, known for its high computational efficiency. This approach, extending to other subsequent calculations, enabled us to numerically solve equations (49) and (53) with a high degree of convergence, without the need to resort to approximations that might otherwise restrict the descriptive potential of the present formulation. Moreover, it is crucial to highlight that when deriving equation (49), which describes the behavior of dynamic solutions corresponding to the second-order derivatives of the scale factor $\eta$, we made no approximations or descriptive impositions. However, when obtaining equation (53), which characterizes the behavior of dynamic solutions corresponding to the first-order derivatives of the scale factor $\eta$, we introduced two impositions: we assumed an explicit time dependence for the variable $u$, allowing the separation of variables $u$ and $t$, and we treated $\dot{u}$ and $u$ as independent variables. Typical solutions of equation (53) are shown in figure 7. When comparing the results of the solutions of equations (49) and (53), it becomes evident that there is a similar behavior in both solutions, at least on a qualitative level. This suggests that the impositions introduced during the development of equation (53) are consistent enough to be applied to the subsequent equations. The first and second-order differential equations for u(t) will prove essential for the determination of the corresponding equations that describe primordial gravitational waves in the realm of a non-commutative formulation. Our results also indicate a strong dependence on the boundary initial conditions applied to the scale factor $u(t)$. For positive $g_{s}$, depending on the boundary initial conditions, our results indicate the presence of a big crunch on the evolution of the Universe, which instead of expanding, violently contracts on itself. For other boundary initial conditions, the Universe experience a deceleration period followed by a moderate acceleration period. For negative $g_{s}$ the results indicate only the deceleration period followed by an moderate acceleration period. Our study reveals the presence of a plethora of solutions in the commutative expansion sector. The solutions are significantly dependent on the choices of parameters, on the running coupling constants, and on boundary conditions.
IV.2 Dynamical equations: Non-commutative approach
@@
Figure 8: Graphical illustrations of the potential $V(\eta,t)$ from equation (63) and corresponding contour plots. In the left pair of figures, the values of the running coupling constants and the parameters of the non-commutative algebra are as follows: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\alpha=1/2$; ${\cal C}=1$. In the right pair of figures, the values of the running coupling constants and the parameters of the non-commutative algebra are as follows: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\alpha=1/2$; ${\cal C}=1$.
From equations (20) and (21), the starting point of our approach is based on the super-Hamiltonian which reads, in the non-commutative BCG, as:
${\cal H}^{NC}=\frac{1}{2}N\Biggl{[}-\frac{1}{\eta}p^{2}_{\eta}-\frac{\gamma}{\eta^{3\alpha}}p_{\eta}+\frac{g_{r}}{\eta}-g_{m}-g_{k}\eta-g_{q}\eta^{2}+g_{\Lambda}\eta^{3}+\frac{g_{s}}{\eta^{3}}+\frac{\alpha}{\eta^{3\alpha-1}}-\frac{\alpha\xi}{\eta^{3\alpha}}+\frac{1}{\eta^{3\alpha}}p_{\xi}\Biggr{]}\,.$ (55)
The Hamilton equation (55) characterizes the time dynamical evolution of the branch-cut Universe parameterized by a dimensionless cosmic scale factor represented, in the present case in which we consider a non-commutative algebraic formalism, by the variable $\eta(t)$. As a peculiarity, the results referring to commutative algebra indicate a late time moderate expansion of the wave function of the Universe and the scale factor.
IV.2.1 Second-order dynamical equations: non-commutative formulation

### 03 Feb 2024 v2

@@ 4 Dynamical Equations
4.1 Dynamical equations: commutative approach
Going a few steps back, the Hořava-Lifshitz commutative BCG action (3.1) is related to the Lagrangian density ${\cal L}_{HL}$, expressed in terms of the standard variable $u(t)$ in the form [8] (see also refs. [38, 50, 49])
@@ 4.1 Dynamical equations: commutative approach
$p_{u}=\frac{\partial{\cal L}_{HL}}{\partial\dot{u}}=-\frac{u\dot{u}}{N}\,.$ (4.3)
@@
${\cal H}_{HL}=\frac{1}{2}\frac{N}{u}\Biggl{[}-p^{2}_{u}+g_{r}-g_{m}u-g_{k}u^{2}-g_{q}u^{3}+g_{\Lambda}u^{4}+\frac{g_{s}}{u^{2}}\Biggr{]}\,.$ (3.6 revisited)
Figure 4: Graphical illustrations of the potential $V(u)$ of equation (4.11), emphasizing the contour lines. The values of the running coupling constants and parameters are, in the left figures, $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$. In the right figures, the values are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\omega=-1$ ${\cal C}=1$.
@@
${\cal H}_{fp}=\frac{1}{2}N\frac{p_{v}}{u^{3\omega}}\,,\quad\mbox{with}\quad p_{v}=-\frac{v\dot{v}}{N}\,,$ (4.5)
resulting, from equation (3.6), assuming the gauge $N=1$, and simplifying the notation, eliminating the HL label, in the super-Hamiltonian:
=12[${\cal H}\!=\!\frac{1}{2}\Biggl{[}-\frac{1}{u}p^{2}_{u}+\frac{g_{r}}{u}-g_{m}-g_{k}u-g_{q}u^{2}+g_{\Lambda}u^{3}+\frac{g_{s}}{u^{3}}+\frac{1}{u^{3\omega}}p_{v}\Biggr{]}.$ (4.6)
From equation (4.6) the following Hamilton equations result:
@@
$\dot{p}_{v}$ $=$ $-\frac{\partial{\cal H}}{\partial v}=0\rightarrow p_{v}={\cal C}\,,$ (4.7)
@@
$\dot{p}_{u}=-\frac{\partial{\cal H}}{\partial u}=-\frac{1}{2u}\Biggl{[}\frac{p^{2}_{u}}{u}-\frac{g_{r}}{u}-g_{k}u-2g_{q}u^{2}-3g_{\Lambda}u^{3}-3\frac{g_{s}}{u^{3}}-3\omega\frac{{\cal C}}{u^{3\omega}}\Biggr{]}\,,$ (4.8)
where ${\cal C}$ is a constant in the time sector. It is important to point out that the relative sign between the term labeled by $g_{\Lambda}$ has changed, in the derivative procedure, in comparison with the $g_{r}$ and $g_{s}$ contributions (see equation (4.6)).
4.1.1 Second-order dynamical equations: commutative formulation
@@
$p_{u}=-u\dot{u}\,,\quad\mbox{and}\quad\dot{p}_{u}=-\dot{u}^{2}-u\ddot{u}\,,$ (4.9)
we obtain from (4.8) the following dynamical equation for the scale factor $u$:
$2u\ddot{u}+\dot{u}^{2}+V(u)=0\,.$ (4.10)
@@
$V(u)=g_{k}+2g_{q}u-3g_{\Lambda}u^{2}+\frac{g_{r}}{u^{2}}+3\frac{g_{s}}{u^{4}}+3\omega\frac{{\cal C}}{u^{3\omega+1}}\,.$ (4.11)
Figure 4 show typical graphical illustrations of the potential $V(u)$, eq. (4.11). The main difference between the curves is related to the sign of the stiff matter running coupling constant, for positive signal it gives rise to a repulsive potential (upper curves) and for negative values an attractive potential around a singularity in the origin.
Figure 5: Sample solutions for the scale factor $u(t)$ given by equation (4.10) for initial conditions $u(0)=0.15$, 0.35, and 1.0. The values of the running coupling constants and parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$.
Figures 5 show typical sample solutions of the second-order time derivative of the scale factor $u(t)$ given in equation (4.10). For all values of the running coupling constants, the results indicate, differently from generally assumed, that the Universe does not originate from a singularity or out of nothing as originally proposed by Alan Guth [6, 7] in the inflation model (see also ref. [65]). Our results are more in tune with a transition phase between the contraction and expansion topological boundaries intermediated by a Riemannian foliated transition region as proposed by the branch-cut gravity.
@@
Figure 6: Graphical illustrations of the potential $V(u,t)$ and the scale factor of equations (4.14) and (4.15) for the commutative approach. In the left-pair of figures, the values of the running coupling constants and parameters are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$. In the right-pair of figures, the values are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\omega=-1$ ${\cal C}=1$.
4.1.2 First-order dynamical equations: commutative formulation
In what follows, we use Cauchy’s implicit function theorem, assuming that the variable $u$ has an explicit time dependence, so the variables $u$ and $t$ can be separated, thus ensuring that these variables are differentiable. We also assume, for simplicity, in view of the high difficulty to solve the above equations, that $\dot{u}$ and $u$ act as independent variables, so from equation (4.10) we have
$2u\frac{d\dot{u}}{dt}+\dot{u}\frac{du}{dt}=-V(u)\rightarrow 2ud\dot{u}+\dot{u}du=-V(u)dt\,.$ (4.12)
@@ 4.1.2 First-order dynamical equations: commutative formulation
$V(u,t)\equiv\frac{1}{3}\Biggl{(}\frac{g_{k}}{u}+2g_{q}-3g_{\Lambda}u+\frac{g_{r}}{u^{3}}+3\frac{g_{s}}{u^{5}}+3\omega\frac{{\cal C}}{u^{3\omega+2}}\Biggr{)}t\,.$ (4.15)
@@
Figure 6 presents graphical representations of equation (4.15). In these contour plots, varying colors indicate the behavior of the potential $V(u,t)$ in regions of dominance, either repulsion or attraction. Lighter colors signify higher intensity or amplitude, while darker colors indicate the opposite. The blue color signifies the region with lower values, typically surrounding the singularity or the repulsive part of the potential, while the cream color designates regions with higher values, where either the singularity or the repulsion of the potential predominates. The white regions, albeit not always clearly visible in some of the figures, corresponds to the exact region of the singularity or the maximum of repulsion. The results regarding the potential $V(u,t)$ are characterized by symmetric distributions of color compositions, indicating that the commutative algebraic structure contributes in a moderate way to the configuration and distribution of matter in the early Universe.
@@
Figure 7 displays typical sample solutions of equation (4.14) for various sets of initial conditions. The outcomes obtained from the dynamic equations suggest that, contingent upon the initial conditions, phases that lead to the big crunch or moderately accelerated phases of the early Universe tend to prevail.
@@
Figure 7: Sample solutions of the first-order time derivative of $u(t)$ as given in equation (4.14) for the commutative formulation. The initial condition is $u(0)$. The values of the running coupling constants and parameters are as follows. Left figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\omega=-1$; ${\cal C}=1$; Middle figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\omega=-1$; ${\cal C}=1$; Right figure: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=-0.4$; $g_{s}=-0.03$; $\omega=-1$; ${\cal C}=1$.
@@
The consistency of the assumptions that led to equation (4.14) and to the potential $V(u,t)$ given by equation (4.15) will be tested in the following by comparing the results that describe the behavior of the solutions of the dynamic equations (4.10) and (4.14). It is important to emphasize that articles addressing this topic typically do not provide solutions to the equations describing the behavior of the Universe’s wave function due to the formal challenges associated with these descriptive equations. These approaches are constrained to performing computational or algebraic calculations for dynamic equations, typically relying on approximations that significantly restrict their scope and predictive capabilities. In our case, from a computational perspective, we utilized the Runge-Kutta-Fehlberg method for solving differential equations, known for its high computational efficiency. This approach, extending to other subsequent calculations, enabled us to numerically solve equations (4.10) and (4.14) with a high degree of convergence, without the need to resort to approximations that might otherwise restrict the descriptive potential of the present formulation. Moreover, it is crucial to highlight that when deriving equation (4.10), which describes the behavior of dynamic solutions corresponding to the second-order derivatives of the scale factor $\eta$, we made no approximations or descriptive impositions. However, when obtaining equation (4.14), which characterizes the behavior of dynamic solutions corresponding to the first-order derivatives of the scale factor $\eta$, we introduced two impositions: we assumed an explicit time dependence for the variable $u$, allowing the separation of variables $u$ and $t$, and we treated $\dot{u}$ and $u$ as independent variables. Typical solutions of equation (4.14) are shown in figure 7. When comparing the results of the solutions of equations (4.10) and (4.14), it becomes evident that there is a similar behavior in both solutions, at least on a qualitative level. This suggests that the impositions introduced during the development of equation (4.14) are consistent enough to be applied to the subsequent equations. The first and second-order differential equations for u(t) will prove essential for the determination of the corresponding equations that describe primordial gravitational waves in the realm of a non-commutative formulation. Our results also indicate a strong dependence on the boundary initial conditions applied to the scale factor $u(t)$. For positive $g_{s}$, depending on the boundary initial conditions, our results indicate the presence of a big crunch on the evolution of the Universe, which instead of expanding, violently contracts on itself. For other boundary initial conditions, the Universe experience a deceleration period followed by a moderate acceleration period. For negative $g_{s}$ the results indicate only the deceleration period followed by an moderate acceleration period. Our study reveals the presence of a plethora of solutions in the commutative expansion sector. The solutions are significantly dependent on the choices of parameters, on the running coupling constants, and on boundary conditions.
4.2 Dynamical equations: Non-commutative approach
@@
Figure 8: Graphical illustrations of the potential $V(\eta,t)$ from equation (4.24) and corresponding contour plots. In the left-pair of figures, the values of the running coupling constants and the parameters of the non-commutative algebra are as follows: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\alpha=1/2$; ${\cal C}=1$. In the right-pair of figures, the values of the running coupling constants and the parameters of the non-commutative algebra are as follows: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=0.03$; $\alpha=1/2$; ${\cal C}=1$.
From equations (3.19) and (3.20), the starting point of our approach is based on the super-Hamiltonian which reads, in the non-commutative BCG, as:
${\cal H}^{NC}=\frac{1}{2}N\Biggl{[}-\frac{1}{\eta}p^{2}_{\eta}-\frac{\gamma}{\eta^{3\alpha}}p_{\eta}+\frac{g_{r}}{\eta}-g_{m}-g_{k}\eta-g_{q}\eta^{2}+g_{\Lambda}\eta^{3}+\frac{g_{s}}{\eta^{3}}+\frac{\alpha}{\eta^{3\alpha-1}}-\frac{\alpha\xi}{\eta^{3\alpha}}+\frac{1}{\eta^{3\alpha}}p_{\xi}\Biggr{]}\,.$ (4.16)
The Hamilton equation (4.16) characterizes the time dynamical evolution of the branch-cut Universe parameterized by a dimensionless cosmic scale factor represented, in the present case in which we consider a non-commutative algebraic formalism, by the variable $\eta(t)$. As a peculiarity, the results referring to commutative algebra indicate a late time moderate expansion of the wave function of the Universe and the scale factor.
4.2.1 Second-order dynamical equations: non-commutative formulation